OpenStudy (anonymous):

determine if (1/3,-5) is a solution to -3x-10y=51

OpenStudy (anonymous):

-3(1/3) -10(-5) = 51 -1 + 50 = 51 49=51 What do you think the answer is?

OpenStudy (anonymous):

2